Determine the operation tables for group G with orders 1, 2, 3 and 4 using the elements a, b, c, and

e as the identity element in an appropriate way.


1
Expert's answer
2021-01-28T15:34:03-0500
  1. There is only one group of order 1, as it consists only of identity element "e". There is no need to do a multiplication table in this case.
  2. For a group of order 2 we have two elements - "e, a". As in a group any element should have an inverse, we will forcefully have "a\cdot a = e" (as "a\cdot e = e\cdot a = a", so the identity element can not be inverse of "a")
  3. For a group of order 3 we also have only one possible group. We have elements "e,a,b". We will study the possible values of "a \cdot b" : it can not be "a" or "b" as in this case we would have "a\cdot b = a \Rightarrow b=e" (or "a\cdot b = b \Rightarrow a=e"), therefore we have "a\cdot b = e". This also means that "a\cdot a \neq e" as the inverse is unique and thus we have "a\cdot a = b". The same reasoning gives us "b\cdot b = a".
  4. For a group of order 4 we, actually, have two possibilities. Let's start by studying the value of "a\cdot a". If "a \cdot a = b", then we forcefully have "a\cdot b = c". Indeed, we can not have "a\cdot b = a" or "=b" and if "a\cdot b=e", then "a\cdot c" can not be neither "e" (as inverse of "a" is "b"), not "a,c" and it can not be "b", as "b = a\cdot a". Therefore we have a group "e, a, a^2,a^3" with "a^4 =e" and the multiplication table becomes trivial. The cases "a\cdot a =c", "b\cdot b =a", "b\cdot b =c", "c\cdot c =a", "c\cdot c =b" are completely analogous as it is just enough to rename the letters a,b and c. Therefore we now need to treat the case when "a^2 = b^2 =c^2=e". But in this case the multiplication table trivially becomes "b\cdot a = a\cdot b =c", "c\cdot a = a\cdot c =b", "b\cdot c = c\cdot b =a".